OpenStudy (anonymous):

Please Help, Create a rational function with a linear binomial in both the numerator and denominator. Part 1. Graph your function using technology. Include the horizontal and vertical asymptotes and the x- and y-intercepts on your graph. Label the asymptotes and intercepts. Part 2. Show all work to identify the vertical asymptote, the x-intercepts, and the y-intercept.

OpenStudy (anonymous):

Rational function looks like a fraction. Linear binomial is something like (x-2). You want similar terms in both the numerator and the denominator, so it might look like\[f(x)=\frac{x-a}{x-b}\]Pick numbers for a and b, and you are on your way.

OpenStudy (anonymous):

does it matter which numbers I pick

OpenStudy (anonymous):

Nope, as long as they aren't zero, or the same as one another.

OpenStudy (anonymous):

ok so a=3 b=5

OpenStudy (anonymous):

I recommend easy numbers: small integers.

OpenStudy (anonymous):

Yeah, that will work.

OpenStudy (anonymous):

ok so I would just graph those and then how would I figure out part 2

OpenStudy (anonymous):

Well, the vertical asymptote is at the undefined value of x. Can you tell me what that is?

OpenStudy (anonymous):

im not really sure im like confused

OpenStudy (anonymous):

Well, is there any number that is excluded from the domain?

OpenStudy (anonymous):

We aren't allowed to divide by zero, so any number that makes the denominator of the function zero can't be in the domain. What value of x would that be?

OpenStudy (anonymous):

um im not really sure would it be like 5

OpenStudy (anonymous):

Exactly. That means x=5 is the vertical asymptote. Now, how do we find the x intercept?

OpenStudy (anonymous):

If the function is intercepting the x axis, what is the y value?

OpenStudy (anonymous):

ok well I graphed \[f(x)=\frac{ x-3 }{ x-5 }\] and got 3 as the x intercept

OpenStudy (anonymous):

Correct. To do it algebraically, you set y (keeping in mind that y = f(x)) to be zero. That will allow you to get that solution. Good. Two down, one to go. How do we find the y intercept?

OpenStudy (anonymous):

do you set x=0

OpenStudy (anonymous):

Yep.

OpenStudy (anonymous):

so would the y intercept be 0.6

OpenStudy (anonymous):

Looks like you've got it. Nice work. Do math every day.

OpenStudy (anonymous):

lol thank you so much this really helped me a lot

OpenStudy (anonymous):

No sweat. Do math every day.